OpenStudy (anonymous):

At Central Park Zoo, the ratio of sea lions to penguins is 4:11. If there are 84 more penguins that sea lions, how many sea lions are there?

hero (hero):

4/11 = x/(x+84) 11x = 4(x+84) 11x = 4x + 336 7x = 336 x = 336/7 x =48 x +84 = 132

hero (hero):

There are 48 sea lions